### v2

# Axis Decomposition for ODRL: Resolving Dimensional Ambiguity in Policy Constraints through Interval Semantics

Abstract
Every ODRL 2.2 constraint compares a single scalar value: (leftOperand, operator, rightOperand). Five of ODRL’s left operands, however, denote multi-dimensional quantities—image dimensions, canvas positions, geographic coordinates—whose specification text explicitly references multiple axes. For these operands, a single scalar constraint admits one interpretation per axis, making policy evaluation non-deterministic. We classify ODRL’s left operands by value-domain structure (scalar, dimensional, concept-valued), grounded in the ODRL 2.2 specification text, and show that dimensional ambiguity is intrinsic to the constraint syntax. We present an axis-decomposition framework that refines each dimensional operand into axis-specific scalar operands and prove four properties: deterministic interpretation, AABB completeness, sound over-approximation under projection, and conservative extension. Conflict detection operates in two layers: per-axis verdicts are always decidable; box-level verdicts compose through Strong Kleene conjunction into a three-valued logic (Conflict, Compatible, Unknown). For ODRL’s disjunctive (odrl:or) and exclusive-or (odrl:xone) logical constraints, where per-axis decomposition does not apply, the framework encodes coupled multi-axis conjectures directly. We instantiate the framework as the ODRL Spatial Axis Profile—15 axis-specific left operands for the five affected base terms—and evaluate it on 117 benchmark problems spanning nine categories across both TPTP FOF (Vampire) and SMT-LIB (Z3) encodings, achieving full concordance between provers. Benchmark scenarios are inspired by constraints arising in cultural heritage dataspaces such as Datenraum Kultur. All meta-theorems are mechanically verified in Isabelle/HOL.
1 Introduction
The Open Digital Rights Language (ODRL) is the W3C Recommendation for expressing digital policies [7], mandated for policy exchange in European dataspace ecosystems including Gaia-X [5], IDSA, and the Eclipse Dataspace Connector [4]. Every ODRL constraint takes the form (leftOperand, operator, rightOperand) a single scalar comparison. This design works for scalar quantities (counts, dates, percentages) but becomes ambiguous when the left operand denotes a multi-dimensional quantity—image dimensions, canvas positions, or geographic coordinates—because the syntax provides one comparison slot for a value that spans multiple axes.
Five of ODRL’s left operands are dimensional. In a dataspace where hundreds of policies govern asset reuse across organisations, this ambiguity compounds into concrete failure modes: non-deterministic enforcement, where the same constraint yields different verdicts depending on which axis an engine selects; inexpressible bounding boxes, where a policy author cannot conjoin per-axis bounds because ODRL provides one operand for multiple dimensions; undetectable self-contradictions, where a single policy simultaneously requires $\texttt{absoluteSize}\leq 600$ and $\texttt{absoluteSize}\geq 800$ on different axes but no engine can flag the conflict without axis-level reasoning; and silent cross-policy incompatibility, where an upstream licence permits $1200\times 900$ px but a downstream policy restricts to $600\times 400$ px, and the supply chain breaks without diagnosis.
We trace these failures through a scenario from the Datenraum Kultur cultural dataspace [9] (Table 1). The Bayerische Staatsbibliothek (BSB) licenses digitised manuscripts at a maximum edge length of 600 pixels [3]. A Dutch museum requests a $1200\times 400$ crop for an exhibition catalogue, restricted to display within the Netherlands. The policy combines dimensional constraints (image size), concept-valued constraints (spatial region, purpose), and scalar constraints—yet no existing mechanism disambiguates which axis the size bound targets, composes per-axis verdicts, or detects conflicts across these heterogeneous domains.
None of the available remedies suffices. External vocabularies (GeoSPARQL [2], EXIF [1]) can describe axes but cannot inject axis identity into the ODRL constraint syntax. Ontology alignment conflates norms with descriptions (Proposition˜2.5). Unit specification resolves “600 what?” but not “600 along which axis?”
Table 1: Three escalating problems: BSB licenses digitised manuscripts at max 600 px per edge for non-commercial European display; a Dutch museum requests a $1200\times 400$ crop for an exhibition catalogue.
Problem Scenario Without framework With framework
Axis ambiguity
Which axis?
Policy: absoluteSize lteq 600.
Request: $1200\times 400$ px.
Width violates; height satisfies.
Engine must guess an axis.
Verdict is non-deterministic.
Decomposed into width lteq 600, height lteq 600.
Width: Conflict.
Height: Compatible.
Conjunction: Conflict.
Multi-axis
composition
Bounding boxes?
BSB intends $w\leq 600\wedge h\leq 600$.
ODRL has one operand for both axes.
Duplication duplicates ambiguity.
No mechanism to conjoin per-axis bounds.
AABB policies inexpressible.
Axis-specific operands compose via and. Each targets one axis; conjunction defines the AABB.
Cross-domain
composition
Verdict composition?
Full policy: size $\leq 600$ px and spatial $\in$ Europe and purpose == non-commercial. Dimensional and KB verdicts have no common framework. Per-operand verdicts compose via three-valued conjunction. One Conflict $\Rightarrow$ overall Conflict.
We present an axis-decomposition framework that resolves all four failure modes. Each dimensional left operand is refined into axis-specific scalar operands; each axis-specific constraint maps to an interval denotation; and conflict detection reduces to interval intersection, yielding a three-valued verdict (Conflict, Compatible, Unknown) that composes soundly across heterogeneous constraint domains.
Contributions.
Classification of ODRL left operands; proof that five are dimensional with n+2n{+}2 ambiguous interpretations each; proof that KB grounding, alignment, and unit specification each fail to resolve the ambiguity (Section˜2).
Axis-decomposition framework with four verified properties: deterministic interpretation, AABB completeness, projection soundness, and conservative extension. All mechanically verified in Isabelle/HOL (Section˜3).
Strong Kleene conflict detection composing per-axis verdicts (always decidable) into box-level and cross-domain verdicts for all three ODRL connectives (and, or, xone), with soundness guarantees under incomplete information (Sections˜3.2 and 3.3).
ODRL Spatial Axis Profile: 15 axis-specific left operands, OWL vocabulary, SHACL validation, and TPTP axiom files (Footnote˜2).
117 dual-encoded benchmarks (TPTP FOF + SMT-LIB) in nine categories, with 100% Vampire/Z3 concordance (Section˜3.6).
2 The Dimensionality Problem
This section formalises the ambiguity illustrated in Table 1 and proves it intrinsic to ODRL’s constraint syntax.
[…]
Definition 2.1 (Constraint).
A constraint is a tuple $c=(\ell,\bowtie,v)$ where $\ell\in\mathcal{L}$ is a left operand drawn from the ODRL vocabulary, $\bowtie\in\mathcal{O}$ is a comparison operator, and $v\in\mathcal{V}$ is a right operand value.
ODRL defines twelve operators partitioned into two groups: six comparison operators $\{\texttt{eq},\texttt{neq},\texttt{lt},\texttt{lteq},\texttt{gt},\texttt{gteq}\}$ that impose numeric or lexicographic order, and six set-based operators (isA, isPartOf, etc.) that test membership. Since set-based operators do not compare along a numeric axis, they are not subject to dimensional ambiguity.
2.1 Dimensional Left Operands
Definition 2.2 (Dimensional Left Operand).
A left operand $\ell$ is dimensional if its value space $\mathcal{V}_{\ell}=D_{1}\times\cdots\times D_{n}$, $n\geq 2$, where each $D_{k}$ is totally ordered.
Five of ODRL’s left operands are dimensional (Table 2). [^1] The count corresponds to the left operands enumerated in the ODRL Vocabulary & Expression 2.2 [6], § 3.
Proposition 2.3 (Dimensional Insufficiency).
For a dimensional left operand $\ell$ with $\mathcal{V}_{\ell}=D_{1}\times\cdots\times D_{n}$, the constraint $(\ell,\bowtie,v)$ with $v\in\mathbb{R}$ admits $n+2$ distinct satisfaction conditions: $n$ axis-specific conditions $\pi_{k}(\ell(w))\bowtie v$ for $1\leq k\leq n$, plus two aggregate conditions $\max_{k}\pi_{k}(\ell(w))\bowtie v$ and $\min_{k}\pi_{k}(\ell(w))\bowtie v$. The constraint syntax specifies neither the axis nor the aggregation mode.
Proof.
Each projection $\pi_{k}$ yields a distinct condition. The aggregate conditions $\max_{k}$ and $\min_{k}$ are distinct from each other and from any single projection when $n\geq 2$: consider $\ell(w)=(v{-}1,v{+}1)$ in $D_{1}\times D_{2}$. Then $\pi_{1}(\ell(w))\bowtie v$ and $\pi_{2}(\ell(w))\bowtie v$ yield different truth values for strict operators, and $\max\bowtie v$ differs from $\min\bowtie v$. ∎
Table 2: Dimensional left operands in ODRL 2.2 and their normative specification text evidencing multi-axis value spaces. Quoted text is verbatim from the ODRL 2.2 ontology [6].
Left Operand $n$ ODRL 2.2 ontology definition
[…]
absoluteSpatial-
Position
2/3 “The absolute spatial positions of four corners of a rectangle on a 2D-canvas or the eight corners of a cuboid in a 3D-space”
relativeSpatial-
Position
2/3 “The relative spatial positions – expressed as percentages of full values – of four corners of a rectangle on a 2D-canvas or the eight corners of a cuboid in a 3D-space”
spatialCoordinates 2/3 “The coordinates MUST include longitude and latitude, they MAY include altitude”
[…]
2.2 Insufficiency of Knowledge Grounding
Proposition 2.4 (KB Cannot Resolve Left-Operand Ambiguity).
For a dimensional left operand $\ell$, the constraint $(\ell,\bowtie,v)$ is ambiguous regardless of how precisely $v$ is typed.
Proof.
Even with $v$ fully grounded (e.g., $1250\in\texttt{xsd:decimal}$), the constraint admits $n+2$ interpretations (Proposition 2.3). An external ontology may identify axes, but the ODRL evaluator still receives $(\ell,\bowtie,v)$ with no axis identity and no aggregation mode. ∎
Proposition 2.5 (Vocabulary Alignment is Unsound).
Declaring $\ell_{a}\;\texttt{owl:equivalentProperty}\;p$, where $p$ is a descriptive property (e.g., exif:width), licenses unsound inferences.
Proof.
owl:equivalentProperty is symmetric. If a policy states $(\ell_{a},\texttt{lteq},1250)$, the alignment entails $\texttt{exif:width}\leq 1250$ as a descriptive fact. This conflates a norm (the asset must satisfy the bound) with a description (the asset does have that value): a 4000 px image violates the constraint but does not cease to have width 4000. ∎
Remark 2.6 (Unit vs. Axis).
ODRL’s odrl:unit resolves right-operand typing (“1250 what?”). Left-operand ambiguity persists: absoluteSize lteq 1250 px leaves open whether width, height, or depth is constrained.
Corollary 2.7 (Necessity of Axis Decomposition).
Dimensional ambiguity can be resolved only by axis-specific operands within the constraint language. Neither knowledge-base grounding, vocabulary alignment, nor unit specification suffices.
3 Axis Decomposition
Definition 3.1 (Axis Decomposition).
For a dimensional left operand $\ell$ with $\mathcal{V}_{\ell}=D_{1}\times\cdots\times D_{n}$, the axis decomposition is a set $\Delta(\ell)=\{\ell_{a_{1}},\ldots,\ell_{a_{n}}\}$ where each $\ell_{a_{k}}$ has value space $\mathcal{V}_{\ell_{a_{k}}}=D_{k}$ and $\ell_{a_{k}}(w)=\pi_{k}(\ell(w))$.
Each decomposed operand isolates a single axis, enabling unambiguous constraints:
Definition 3.2 (Axis-Specific Constraint).
An axis-specific constraint $c=(\ell_{a_{k}},\bowtie,v)$ with $v\in D_{k}$ is satisfied by execution context $w$ iff:
$w\models c\;\;\iff\;\;\ell_{a_{k}}(w)\!\downarrow\;\land\;\ell_{a_{k}}(w)\;\bowtie\;v$
When $\ell_{a_{k}}(w)$ is undefined, $w\not\models c$ (closed-world per axis: missing data entails non-satisfaction). This is distinct from an unconstrained axis—one for which no constraint targets $a_{k}$—which is handled at the box level (Section˜3.2).
Theorem 3.3 (Semantic Determinism).
Every axis-specific constraint has a unique satisfaction condition.
Proof.
$\ell_{a_{k}}(w)\in D_{k}$ and $v\in D_{k}$ are elements of a totally ordered domain. The truth value of $\ell_{a_{k}}(w)\bowtie v$ is uniquely determined. ∎
Remark 3.4 (Aggregation via Composition).
The aggregate conditions from Proposition 2.3 reduce to standard ODRL composition over axis-specific constraints: $\max_{k}\pi_{k}(\ell(w))\leq v$ is equivalent to conjunction (all axes satisfy), while $\max_{k}\pi_{k}(\ell(w))\geq v$ is equivalent to disjunction (at least one axis satisfies). Both are expressible using and/or over operands from $\Delta(\ell)$. The axis decomposition thus resolves both the axis-selection ambiguity and the aggregation-mode ambiguity.
3.1 Interval Denotation
[…]
Definition 3.5 (Interval Denotation).
For $c=(\ell_{a_{k}},\bowtie,v)$ with $v\in D_{k}$ and $\bowtie\in\mathcal{O}_{\Delta}$:
$\llbracket\ell_{a_{k}}\texttt{ eq }v\rrbracket$ $=[v,v]$ (1)
$\llbracket\ell_{a_{k}}\texttt{ lteq }v\rrbracket$ $=[\inf D_{k},\;v]$ (2)
$\llbracket\ell_{a_{k}}\texttt{ gteq }v\rrbracket$ $=[v,\;\sup D_{k}]$ (3)
$\llbracket\ell_{a_{k}}\texttt{ lt }v\rrbracket$ $=[\inf D_{k},\;v)$ (4)
$\llbracket\ell_{a_{k}}\texttt{ gt }v\rrbracket$ $=(v,\;\sup D_{k}]$ (5)
Definition 3.6 (Box Denotation).
For a conjunction $C$ of axis-specific constraints over $\Delta(\ell)=\{\ell_{a_{1}},\ldots,\ell_{a_{n}}\}$:
$\llbracket C\rrbracket=I_{1}\times\cdots\times I_{n}\quad\text{where}\quad I_{k}=\bigcap\{\llbracket c_{j}\rrbracket\mid c_{j}\text{ targets }a_{k}\}$
If no constraint targets $a_{k}$, then $I_{k}=D_{k}$.
Theorem 3.7 (AABB Completeness).
  1. 1.
    Every conjunction of axis-specific constraints defines an axis-aligned bounding box $R=I_{1}\times\cdots\times I_{n}$.
  2. 2.
    Every AABB in $D_{1}\times\cdots\times D_{n}$ is expressible as such a conjunction.
Proof.
(1) By Definition 3.6. (2) Each face $l_{k},u_{k}]$ is expressed as $(\ell_{a_{k}},\texttt{gteq},l_{k})\;\texttt{and}\;(\ell_{a_{k}},\texttt{lteq},u_{k})$. ∎
Theorem 3.8 (Projection Soundness).
Let $\llbracket C\rrbracket=I_{1}\times\cdots\times I_{n}$. Then $\mathbf{v}\in\llbracket C\rrbracket$ iff $\pi_{k}(\mathbf{v})\in I_{k}$ for all $k$.
Proof.
Immediate from the definition of Cartesian product. ∎
Remark 3.9 (Over-Approximation).
For non-axis-aligned regions, only the forward direction holds: the AABB is a sound over-approximation.
3.2 Conflict Detection
Conflict detection operates in two layers: per-axis verdicts are always definite; box-level verdicts introduce Unknown when an axis is unconstrained.
3.2.1 Per-Axis Verdicts
Definition 3.10 (Per-Axis Verdict).
For axis-specific constraints $c_{1},c_{2}$ over the same axis $a_{k}$:
$\operatorname{verdict}_{a_{k}}(c_{1},c_{2})=\begin{cases}\textsc{Conflict}&\text{if }\llbracket c_{1}\rrbracket\cap\llbracket c_{2}\rrbracket=\emptyset\\ \textsc{Compatible}&\text{if }\llbracket c_{1}\rrbracket\cap\llbracket c_{2}\rrbracket\neq\emptyset\end{cases}$ (6)
Definition 3.11 (Per-Axis Subsumption).
For axis-specific constraints $c_{1},c_{2}$ over the same axis $a_{k}$:
$\operatorname{subsumes}_{a_{k}}(c_{1},c_{2})=\begin{cases}\textsc{Confirmed}&\text{if }\llbracket c_{1}\rrbracket\subseteq\llbracket c_{2}\rrbracket\\ \textsc{Refuted}&\text{if }\llbracket c_{1}\rrbracket\not\subseteq\llbracket c_{2}\rrbracket\end{cases}$ (7)
Lemma 3.12 (Conflict Propagation).
If $\operatorname{subsumes}_{a_{k}}(c_{1},c_{2})=\textsc{Confirmed}$ and $\operatorname{verdict}_{a_{k}}(c_{2},c_{3})=\textsc{Conflict}$, then $\operatorname{verdict}_{a_{k}}(c_{1},c_{3})=\textsc{Conflict}$.
Proof.
$\llbracket c_{1}\rrbracket\subseteq\llbracket c_{2}\rrbracket$ and $\llbracket c_{2}\rrbracket\cap\llbracket c_{3}\rrbracket=\emptyset$ imply $\llbracket c_{1}\rrbracket\cap\llbracket c_{3}\rrbracket=\emptyset$. ∎
3.2.2 Box-Level Verdicts
Definition 3.13 (Box Verdict).
Given constraint sets $C_{1},C_{2}$ over $\Delta(\ell)=\{\ell_{a_{1}},\ldots,\ell_{a_{n}}\}$, the per-axis status is:
$V_{k}=\begin{cases}\operatorname{verdict}_{a_{k}}(c_{1}^{(k)},c_{2}^{(k)})&\text{if both }C_{1},C_{2}\text{ constrain }a_{k}\\ \textsc{Unknown}&\text{otherwise}\end{cases}$ (8)
The box verdict aggregates:
$\operatorname{boxVerdict}(C_{1},C_{2})=\begin{cases}\textsc{Conflict}&\text{if }\exists k\colon V_{k}=\textsc{Conflict}\\ \textsc{Compatible}&\text{if }\forall k\colon V_{k}=\textsc{Compatible}\\ \textsc{Unknown}&\text{otherwise}\end{cases}$ (9)
Table 3: Verdict correspondence across reasoning layers. Per-axis verdicts map to prover outputs; Unknown arises at the box level when an axis lacks constraints and is not submitted to the prover.
Relation Verdict Encoding SZS Z3
Conflict Conflict $\llbracket c_{1}\rrbracket\cap\llbracket c_{2}\rrbracket=\emptyset$ Theorem unsat
Compatible $\llbracket c_{1}\rrbracket\cap\llbracket c_{2}\rrbracket\neq\emptyset$ Theorem sat
[…]
Refuted $\llbracket c_{1}\rrbracket\not\subseteq\llbracket c_{2}\rrbracket$ CounterSat sat
Unknown (either relation) (axis unconstrained) — not submitted —
[…]
Definition 3.14 (Box Subsumption).
Given constraint sets $C_{1},C_{2}$ over $\Delta(\ell)$, the per-axis subsumption status is determined by Definition 3.11:
$S_{k}=\begin{cases}\operatorname{subsumes}_{a_{k}}(c_{1}^{(k)},c_{2}^{(k)})&\text{if both }C_{1},C_{2}\text{ constrain }a_{k}\\ \textsc{Unknown}&\text{otherwise}\end{cases}$ (10)
$\operatorname{boxSubsumes}(C_{1},C_{2})=\begin{cases}\textsc{Confirmed}&\text{if }\forall k\colon S_{k}=\textsc{Confirmed}\\ \textsc{Refuted}&\text{if }\exists k\colon S_{k}=\textsc{Refuted}\\ \textsc{Unknown}&\text{otherwise}\end{cases}$ (11)
Remark 3.15 (Choice of Three-Valued Semantics).
The verdict algebra instantiates strong Kleene three-valued logic ($K_{3}$): conjunction is $\min$ and disjunction is $\max$ over $F<\bot<T$, where $\textsc{Conflict}\mapsto F$, $\textsc{Unknown}\mapsto\bot$, $\textsc{Compatible}\mapsto T$. Heyting semantics ($\neg\bot=F$) is unsound: it would treat an unresolved branch as provably non-overlapping, fabricating exclusivity in xone. Bochvar semantics ($T\cup\bot=\bot$) is too conservative: a single definite branch cannot resolve disjunction. Strong Kleene uniquely satisfies both requirements: $T\cup\bot=T$ (disjunction resolves) and $\neg\bot=\bot$ (exclusivity demands proof).
Proposition 3.16 (No False Verdicts Under Incomplete Information).
If $\operatorname{boxVerdict}(C_{1},C_{2})=\textsc{Conflict}$, the verdict holds for every completion of the unconstrained axes. If $\operatorname{boxVerdict}(C_{1},C_{2})=\textsc{Compatible}$, every axis has been examined and admits overlap. Unknown is returned if and only if no axis proves conflict and at least one axis lacks a definite verdict.
Definition 3.17 (Request Satisfaction).
For request $\mathbf{v}=(v_{1},\ldots,v_{n})$ and constraint set $C$:
$\operatorname{satisfied}(\mathbf{v},C)=\begin{cases}\textsc{Yes}&\text{if }\forall k:v_{k}\in I_{k}\\ \textsc{No}&\text{if }\exists k:v_{k}\notin I_{k}\end{cases}$
Remark 3.18 (Deontic Conflict).
When comparing a permission and prohibition over the same action, deontic conflict occurs when scopes overlap: $\llbracket C_{P}\rrbracket\cap\llbracket C_{Q}\rrbracket\neq\emptyset$. This is the inverse of constraint satisfiability (Definition 3.10): disjoint scopes prevent deontic conflict, while overlapping scopes create it.
3.3 Composition
Within a single policy, and produces box denotations (Definition 3.6), or produces unions of boxes, and xone requires exactly one satisfiable branch.
Cross-policy disjunction.
$\operatorname{verdict}_{\texttt{or}}=\begin{cases}\textsc{Compatible}&\text{if }\exists\,\text{branch pair}\colon\operatorname{boxVerdict}=\textsc{Compatible}\\ \textsc{Conflict}&\text{if }\forall\,\text{branch pairs}\colon\operatorname{boxVerdict}=\textsc{Conflict}\\ \textsc{Unknown}&\text{otherwise}\end{cases}$ (12)
Cross-policy exclusive disjunction.
$\operatorname{verdict}_{\texttt{xone}}=\begin{cases}\textsc{Compatible}&\text{if }\exists!\,\text{branch pair}\colon\operatorname{boxVerdict}=\textsc{Compatible}\\ &\quad\text{and all others }=\textsc{Conflict}\\ \textsc{Conflict}&\text{if }\forall\,\text{branch pairs}\colon\operatorname{boxVerdict}=\textsc{Conflict}\\ \textsc{Unknown}&\text{otherwise}\end{cases}$ (13)
Theorem 3.19 (Composition Soundness).
If $\operatorname{boxVerdict}(C_{1},C_{2})=\textsc{Conflict}$, then no value in $D_{1}\times\cdots\times D_{n}$ simultaneously satisfies both constraint sets. Analogously for disjunctive and exclusive compositions.
Proof.
Conjunction. $\exists k\colon V_{k}=\textsc{Conflict}$ means $\llbracket c_{1}^{(k)}\rrbracket\cap\llbracket c_{2}^{(k)}\rrbracket=\emptyset$. By Theorem 3.8, membership in a box requires membership in every axis projection; since projection $k$ admits no common point, no element of $D_{1}\times\cdots\times D_{n}$ lies in both boxes. Disjunction. All branch pairs yield Conflict: $\forall i,j\colon\llbracket B_{i}\rrbracket\cap\llbracket B_{j}\rrbracket=\emptyset$. Therefore $\bigcup_{i}\llbracket B_{i}\rrbracket\cap\bigcup_{j}\llbracket B_{j}\rrbracket=\bigcup_{i,j}(\llbracket B_{i}\rrbracket\cap\llbracket B_{j}\rrbracket)=\emptyset$. Exclusive disjunction. Reduces to the disjunction case. ∎
3.3.1 Cross-Domain Composition
A complete policy may combine dimensional, concept-valued, and scalar constraints. These target distinct left operands with disjoint value spaces.
Definition 3.20 (Cross-Domain Verdict).
For per-operand verdicts $V_{1},\ldots,V_{m}$ (from any combination of dimensional, concept-valued, or scalar evaluation):
$\operatorname{verdict}_{\times}(V_{1},\ldots,V_{m})=\begin{cases}\textsc{Conflict}&\text{if }\exists i\colon V_{i}=\textsc{Conflict}\\ \textsc{Compatible}&\text{if }\forall i\colon V_{i}=\textsc{Compatible}\\ \textsc{Unknown}&\text{otherwise}\end{cases}$ (14)
Theorem 3.21 (Cross-Domain Soundness).
If $\operatorname{verdict}_{\times}=\textsc{Conflict}$, then no execution context simultaneously satisfies both policies.
Proof.
Some operand $\ell_{i}$ admits no satisfying value. Since operands from different classes have disjoint value spaces, satisfying other operands cannot provide a witness for $\ell_{i}$. ∎
Running example (Table 1, Problem 3). Per-operand verdicts for the BSB/Dutch museum scenario. The first two verdicts arise from interval comparison (this paper); the last two from KB grounding [8]. The cross-domain verdict composes only the labels:
$V_{\text{width}}$ $=\textsc{Conflict}$ $1200\notin(0,600]$
$V_{\text{height}}$ $=\textsc{Compatible}$ $400\in(0,600]$
$V_{\text{spatial}}$ $=\textsc{Compatible}$ $\text{NL}\preceq\text{Europe}$
$V_{\text{purpose}}$ $=\textsc{Unknown}$ $\text{exhibition}\sqsubseteq\text{NonCommercial?}$
$\operatorname{verdict}_{\times}=\textsc{Conflict}$: the width violation alone blocks the request, regardless of the unresolved purpose question.
3.4 Policy Quality Analysis
The per-axis and box-level definitions support three design-time analyses.
Self-contradiction.
A policy requiring width lteq 600 and width gteq 800 under conjunction: $(0,600]\cap[800,\infty)=\emptyset$, so $\operatorname{boxVerdict}=\textsc{Conflict}$. No image satisfies both.
Redundancy.
width lteq 600 and width lteq 1200 under conjunction: $(0,600]\subseteq(0,1200]$, so $\operatorname{subsumes}_{a_{\text{w}}}(c_{1},c_{2})=\textsc{Confirmed}$. The weaker bound is redundant.
Definition 3.22 (Constraint Completeness).
A constraint set $C$ over $\Delta(\ell)$ is axis-complete if for every $a_{k}\in\{a_{1},\ldots,a_{n}\}$, there exists $c\in C$ targeting $a_{k}$. The coverage of $C$ is $\operatorname{cov}(C)=\{a_{k}\mid\exists c\in C\text{ targeting }a_{k}\}$.
Supply-chain refinement.
Upstream permits $1200\times 900$ px; downstream permits $600\times 400$ px. Both axes subsume: $\operatorname{boxSubsumes}=\textsc{Confirmed}$.
Remark 3.23 (Expressiveness Boundary).
The decomposition is limited to axis-aligned regions. Cross-axis constraints—such as total pixel area ($w\times h\leq A$), aspect ratio ($w/h\in[r_{1},r_{2}]$), or crop containment ($x+w_{\text{crop}}\leq W_{\text{canvas}}$)—are not expressible within ODRL’s scalar constraint pattern, which provides a single left-operand slot per constraint. This is a limitation of the language, not the decomposition. A profile may introduce derived scalar operands (e.g., oax:totalPixels, oax:aspectRatio) that pre-compute cross-axis quantities, at which point they require no decomposition and fall under ODRL’s standard scalar evaluation.
Remark 3.24 (Discrete Domains).
Interval denotation applies to both continuous ($\mathbb{R}$) and discrete ($\mathbb{N}$) domains. For pixel dimensions where $D_{k}=\mathbb{N}_{\geq 0}$, half-open intervals simplify: $(v,\infty)$ over $\mathbb{N}$ equals $v{+}1,\infty)$. Conflict detection and subsumption are unchanged.
Remark 3.25 (Domain Bounds).
The exact domain $D_{k}$ for each axis-specific operand is fixed by the profile instantiation (Footnote˜2). For size operands, $D_{k}=(0,\infty)$ since zero collapses the dimension; for relative positions, $D_{k}=[0,100]$; for geographic coordinates, $D_{k}$ follows WGS 84 bounds. All theorems in this section are parametric in $D_{k}$ and hold for any totally ordered domain—open, closed, bounded, or unbounded.
Remark 3.26 (Profile Independence and Expressiveness Limits).
All definitions and theorems in this section are parametric in the axis domains $D_{1},\ldots,D_{n}$: they hold for any profile that decomposes a dimensional left operand into axes satisfying three conditions:
  1. 1.
    each $D_{k}$ is totally ordered (not partially ordered),
  2. 2.
    each $D_{k}$ is non-wrapping (no identification of endpoints), and
  3. 3.
    axes are independent (satisfaction on one axis does not constrain another).
Four classes of dimensional constraint fall outside this scope: (i) circular domains such as longitude at the antimeridian or rotation angles, where regions may cross the wrap point and require unions of intervals; (ii) coupled axes such as 3D orientation, where axis independence fails; (iii) cross-axis constraints such as area or aspect ratio, which ODRL’s scalar pattern cannot express (Remark 3.23); and (iv) non-numeric structured spaces such as colour gamuts. Of these, only the antimeridian case arises in current ODRL practice; extending the interval algebra to circular domains is future work.
Proposition 3.27 (Conservative Extension).
Adding $\Delta(\ell)$ to the ODRL vocabulary is a conservative extension: every constraint using base operands retains its original semantics.
Proof.
Each $\ell_{a_{k}}\in\Delta(\ell)$ has a fresh IRI. The relation $\ell_{a_{k}}\;\texttt{skos:broader}\;\ell$ records specialisation. No SKOS axiom modifies ODRL constraint evaluation, which dispatches on exact left-operand IRIs. A profile-unaware engine ignores unrecognised IRIs per the ODRL profile mechanism [7]; a profile-aware engine gains per-axis precision without altering existing verdicts. ∎
3.5 Mechanical Verification
All meta-theorems in Sections 33.3.1 have been mechanically verified in Isabelle/HOL 2025. The formalisation comprises four theories (${\sim}600$ lines):
ODRL_Interval Layer 0 Interval denotation
ODRL_Axis Layer 1 Per-axis verdicts
ODRL_Box Layer 2 Box composition
ODRL_CrossDomain Layer 3 Cross-domain aggregation
Verified results: Theorems 3.33.21, Lemma 3.12, and end-to-end bridge theorems connecting verdict-level results to set-theoretic properties. The BSB scenario is included as a concrete regression test. The full proof script builds in under 30 s with no sorry. [^2] Isabelle theories and benchmark suite available upon publication. We instantiate the framework as the ODRL Spatial Axis Profile (OAX): an OWL 2 ontology (http://w3id.org/odrl/spatial-axis#) declaring 15 axis-specific LeftOperand instances with SKOS links to parent operands; a SHACL shapes graph constraining odrl:rightOperand to xsd:decimal with domain-appropriate bounds (size: $(0,\infty)$; latitude: [90,90][-90,90]; longitude: [180,180][-180,180]; relative: $[0,100]$); and TPTP axiom files encoding the interval predicates. The SHACL shapes catch malformed policies at authoring time and ground the domain bounds used in automated reasoning.
TPTP axiom files.
The profile’s formal semantics are encoded in two include files: AXIS000-0.ax (interval predicates and their interaction with the ordering) and ORD001-0.ax (density axioms for open-boundary witnesses). Together they provide 13 axioms that, combined with per-problem ordering chains, suffice for automated conflict detection.
Table 4: The 15 axis-specific left operands of the ODRL Spatial Axis Profile. Each IRI composes the base operand name with the axis suffix. Value domain $D_{k}$ is fixed by the profile.
Base operand Axis Axis operand (IRI) $D_{k}$
absoluteSize width oax:absoluteSizeWidth $(0,\infty)$
height oax:absoluteSizeHeight $(0,\infty)$
depth oax:absoluteSizeDepth $(0,\infty)$
relativeSize width oax:relativeSizeWidth $(0,100]$
height oax:relativeSizeHeight $(0,100]$
depth oax:relativeSizeDepth $(0,100]$
absoluteSpatialPosition x oax:absoluteSpatialPositionX $\mathbb{R}$
y oax:absoluteSpatialPositionY $\mathbb{R}$
z oax:absoluteSpatialPositionZ $\mathbb{R}$
relativeSpatialPosition x oax:relativeSpatialPositionX $[0,100]$
y oax:relativeSpatialPositionY $[0,100]$
z oax:relativeSpatialPositionZ $[0,100]$
spatialCoordinates lon. oax:spatialCoordinatesLongitude [180,180][-180,180]
lat. oax:spatialCoordinatesLatitude [90,90][-90,90]
alt. oax:spatialCoordinatesAltitude $\mathbb{R}$
3.6 Evaluation
We evaluate the framework on 117 synthetic benchmark problems organised in nine categories (Table 5).
Table 5: Benchmark categories. Connective column indicates the ODRL constraint composition tested; implicit AND enables per-axis decomposition, while or/xone couple axes.
Cat. Name # Axes Connective Key test
A SingleAxis 15 1 AND (implicit) All 5 operators, density
B Box2D 11 2 AND Kleene conjunction
C Box3D 12 3 AND Associativity, 3D subsumption
D Composition 12 4 AND Mixed ops, scaling
E CrossDomain 12 2–4 AND Heterogeneous axes
F PolicyQuality 17 1–4 AND Difficulty ladder (3–66 ords)
G Boundary 16 1–3 AND All 10 operator pairs at $v$
H LogicalOr 12 2–3 odrl:or Coupled axes, cross-pick
I LogicalXone 10 2–3 odrl:xone Exclusive-or, L-shaped regions
Total 117
Each problem encodes a pair of ODRL constraints as both a TPTP FOF conjecture (for Vampire) and an SMT-LIB formula (for Z3). Categories A–E test implicit-AND composition at increasing dimensionality (1–4 axes). Category F (PolicyQuality) provides a graduated difficulty ladder from trivial (3 constants, 1 axis) to hard (12 constants, 4 axes, 66 ordering axioms). Category G (Boundary) systematically covers all ten operator pairs at a shared boundary point. Categories H and I test odrl:or and odrl:xone logical constraints, where per-axis decomposition does not apply and the prover must reason over coupled multi-axis disjunctions.
Every benchmark carries an expected verdict (Conflict, Compatible, Subsumes, or Refuted) mapped to both SZS status (Theorem/CounterSatisfiable) and SMT status (unsat/sat). Across all 117 problems, Vampire and Z3 agree on every verdict, yielding 100% cross-prover concordance. This confirms that the axiomatised interval framework (FOF) is faithful to the built-in arithmetic semantics (SMT-LIB).
The benchmark scenarios are inspired by image-resolution and spatial constraints arising in cultural heritage dataspaces such as Datenraum Kultur. The BSB scenario of Table 1 is included as a concrete test case: both provers confirm that the cross-domain verdict is Conflict and that width is the sole conflicting axis.
4 Related Work
A full discussion of related work on ODRL profiles, formal semantics, and spatial constraints in digital rights will appear in the next version.
5 Conclusion
We identified a structural mismatch in ODRL 2.2 between the multi-dimensional semantics of five spatial left operands and the scalar constraint pattern. This mismatch causes semantic ambiguity, expressiveness loss, and unsound reasoning in policy enforcement.
The ODRL Spatial Axis Profile resolves this by introducing 15 axis-specific left operands—three per affected base term—operating entirely within ODRL’s constraint pattern with no syntactic extensions and full backward compatibility. All meta-theorems are mechanically verified in Isabelle/HOL.
Future work includes public release of the profile to the W3C ODRL Community Group, integration with ODRL-SA for design-time conflict detection, and validation on production dataspace deployments.